kalynnnn:

Of A line goes through the points (-2, 4) and (0, -1). (a) What is the slope of the line? Show your work (b) Write the equation of the line in point-slope form. Show your work (c) Write the equation of the line in slope-intercept form. Show your work

Vocaloid:

First thing is to use the slope formula to find m = (y2-y1)/(x2-x1) Point slope form is then (y-y1) = m(x-x1) where (x1,y1) is one of the given points, just pick one After that you will want to distribute the m and re-write in y = mx + b